OpenStudy (texaschic101):

What does it mean when it asks to write the solution set with y arbitrary ?

OpenStudy (texaschic101):

I did a system of equations and found them to be infinite solutions. Now it says write solution set with y arbitrary. What does that mean ?

OpenStudy (texaschic101):

@amistre64 ....please help

OpenStudy (texaschic101):

does it mean set y as the subject and x can be any number being there is infinite solutions ?

OpenStudy (anonymous):

y is arbitrary means you want y to be your independent variable. That means you want an equation with x = some expression involving y. example: x = 2y -3

OpenStudy (texaschic101):

Oh....I had it backwards....set x =.....thank you

OpenStudy (texaschic101):

4x - 5y = 3 -8x + 10y = -6 ------------ infinite solutions 4x - 5y = 3 4x = 5y + 3 x = 5/4y + 3/4 <-- so this is arbitrary ??

OpenStudy (texaschic101):

@thomaster .....is this right ?

OpenStudy (texaschic101):

@TuringTest ...please check this

thomaster (thomaster):

No idea xD I'm bad at math

OpenStudy (texaschic101):

lol.....you could have fooled me, math genius. I got confused on it too

OpenStudy (texaschic101):

@PixieDust1 ...can you help

OpenStudy (texaschic101):

@ash2326 ...can you take a peek at this and tell me if it is correct

OpenStudy (ash2326):

yes, it's correct.

OpenStudy (texaschic101):

would it have mattered if I used the other equation ?

OpenStudy (texaschic101):

-8x + 10y = -6 -8x = 10y - 6 x = -10/8y + 6/8 oh...they are the same once reduced.....so I guess it does not matter

OpenStudy (texaschic101):

thank you ash :)

OpenStudy (ash2326):

Yeah :)